OpenStudy (vampireprincess):

What is the area of the regular hexagon with a side length of 3 cm and an apothem of 2.6 cm? A. 15.6 cm2 B. 23.4 cm2 C. 46.8 cm2 D. 31.2 cm2

OpenStudy (vampireprincess):

this is the pic for this question plz help

OpenStudy (michele_laino):

we have to apply this formula: \[\Large {\text{area = }}\frac{{{\text{perimeter }} \times {\text{apothem}}}}{{\text{2}}}\]

OpenStudy (vampireprincess):

2.6 is the height and 3 is the base so how would you the numbers in your equation

OpenStudy (michele_laino):

hint: perimeter= 6* side= 6*3=...

OpenStudy (michele_laino):

so, perimeter=...?

OpenStudy (michele_laino):

what is: 6*3=...?

OpenStudy (vampireprincess):

18

OpenStudy (michele_laino):

ok! Now the requested area is: \[{\text{area = }}\frac{{{\text{perimeter }} \times {\text{apothem}}}}{{\text{2}}} = \frac{{18 \times 2.6}}{2} = ...\]

OpenStudy (vampireprincess):

18*2.6=46.8 AND THEN divide that with 2 and you get 23.4

OpenStudy (michele_laino):

that's right!

OpenStudy (vampireprincess):

so the answer is 23.4?

OpenStudy (michele_laino):

yes! the area is 23.4 cm^2

OpenStudy (vampireprincess):

can you help me on one more?

OpenStudy (michele_laino):

ok!

OpenStudy (vampireprincess):

What is the area of the figure? A. 279 in2 B. 342 in2 C. 389 in2 D. 398 in2

OpenStudy (michele_laino):

we can view your figure as the union of two rectangles, like this: